The positive geometric mean of two numbers is 8, and the difference of the numbers is 30. Find the numbers.
Answer by Boreal(15235)

the square root (2 numbers so square root) of their product is equal to 8
therefore the product is 64
a*b=64
a-b=30
a=b+30
substitute
(b+30)*b=64
b^2+30b-64=0
(b+32)(b-2)=0
b=-32 and b=2
a=-2 and a=32, want only the positive root.
the numbers are 2 and 32.